OpenStudy (anonymous):

What is the 9th term in this pattern: 2,8,32,128

OpenStudy (anonymous):

I got 131, 072. Is this correct?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

yep, correct! a(n)=2*4^(n-1) a(9)=2*4^(9-1)=131072